OpenStudy (anonymous):

Solve for Theta in the given Interval.. I always get confused in these problems..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\cos (\theta) = -\sqrt{\frac{3}{2}}\]for \[-\pi \leq \theta \leq -\frac{\pi}{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Are you allowed to use a calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Not even a graphing calculator? Not sure how you're expected to do this, then...